### v1

@@ Abstract
We review the literature on boundary discontinuity (BD) designs, a powerful non-experimental research methodology that identifies causal effects by exploiting a thresholding treatment assignment rule based on a bivariate score and a boundary curve. This methodology generalizes standard regression discontinuity designs based on a univariate score and scalar cutoff, and has specific challenges and features related to its multi-dimensional nature. We synthesize the empirical literature by systematically reviewing over $80$ empirical papers, tracing the method’s application from its formative uses to its implementation in modern research. In addition to the empirical survey, we overview the latest methodological results on identification, estimation and inference for the analysis of BD designs, and offer recommendations for practice.
Keywords: regression discontinuity, treatment effects estimation, causal inference.
@@ 1 Introduction
The Regression Discontinuity (RD) design is one of the leading observational methods for program evaluation and causal inference (see Abadie-Cattaneo_2018_ARE; Hernan-Robins_2020_Book, for overviews and further references). In its canonical form, a binary treatment is assigned to units whose value of a score is equal to or above a known cutoff, and not assigned to units whose score is below the cutoff. Under the key assumption that all observable and unobservable pretreatment characteristics vary smoothly around the cutoff, the discontinuous change generated by the hard-thresholding treatment assignment rule based on the univariate score and scalar cutoff can be used to learn about causal treatment effects. Lee-Lemieux_2010_JEL offer an early review with a list of standard RD empirical applications, Cattaneo-Titiunik_2022_ARE give a recent overview of the RD methodological literature, and Cattaneo-Idrobo-Titiunik_2020_CUP; Cattaneo-Idrobo-Titiunik_2024_CUP provide a practical introduction to modern RD methods and practice.
@@
The Boundary Discontinuity (BD) design generalizes the RD design to the case where the score is bivariate, and the treatment is assigned to units according to the location of their score relative to a known boundary that splits its support into two disjoint regions. This setup is sometimes called the Multi-Score RD design (Papay-Willett-Murnane_2011_JoE; Reardon-Robinson_2012_JREE; Wong-Steiner-Cook_2013_JEBS), or the Geographic RD design (Keele-Titiunik_2015_PA; Keele-Titiunik-Zubizarreta_2015_JRSSA; Keele-Titiunik_2016_PSRM; Keele-etal_2017_AIE; Galiani-McEwan-Quistorff_2017_AIE; Rischard-Branson-Miratrix-Bornn_2021_JASA; Diaz-Zubizarreta_2023_AOAS). In this chapter, we discuss the empirical strategies most commonly employed for the analysis and interpretation of BD designs, and review the most recent theoretical and methodological results characterizing their econometric properties (Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location; Cattaneo-Titiunik-Yu_2025_BDD-Pooling).
@@
We introduce the BD design formally in Section 2, and then outline our review of the empirical literature in Economics, Political Science, Education, and other disciplines. Table 5 reports over $80$ empirical papers, which form the basis for the organization of this chapter. We find that the overwhelming majority of papers collapse the bivariate score into a univariate distance score measuring the shortest distance to the treatment assignment boundary, and then report a single average treatment effect estimated by pooling all the observations that are close to the boundary, regardless of their specific location. In contrast, only a handful of empirical papers investigate heterogeneity by localizing to specific regions or points along the boundary, despite the potential importance of this rich information for policy evaluation and decision-making.
@@
Following our review of the empirical literature, in Section 3 we discuss pooling-based methods. We begin by discussing the seminal work of Card-Krueger_1994_AER, who appear to be the first to employ the BD design conceptually, although they did not localize to the assignment boundary for estimation purposes. We then discuss holmes1998effect and black1999better, who appear to be the first two papers to employ localization to the assignment boundary explicitly as their main empirical strategy. Subsequent empirical work has recognized the importance of localization to the boundary, but has considered different empirical strategies for treatment effect estimation via local flexible regression methods. One of the most influential papers is dell2010persistent, who incorporated a polynomial expansion of the bivariate location score as part of the local regression specification. More recently, an alternative empirical approach views the estimation based on distance to the boundary as a pooled univariate RD design, and thus estimates treatment effects using a local regression including a polynomial expansion of the univariate distance score. Recent examples include ito2020willingness and dehdari2022origins. Many empirical papers also include boundary-segment fixed effects in their local regression specifications, and some include interactions with the treatment indicator and/or between the polynomial expansions of the univariate distance score and the bivariate location score. Furthermore, for simplicity, or because distinct treatments occur in different regions of the boundary, some papers reduce the BD design to boundary-segment-specific univariate RD designs—see ou2010leave, LondonoVelezRodriguezSanchez_2020_AEJ and salti2022impact. Putting all together, Section 3 discusses eight distinct pooled local regression specifications commonly found in empirical work.
@@
Building on our review of empirical strategies used for pooled estimation in BD designs, in Section 3.1 we also overview the main methodological results in Cattaneo-Titiunik-Yu_2025_BDD-Pooling for pooling-based estimation and inference. These methods are viewed as univariate RD designs based on the shortest distance to the assignment boundary, and hence are related to identification, estimation, and inference over a specific one-dimensional submanifold curve on the plane, defined by the assignment boundary. Recognizing this mathematical structure, Cattaneo-Titiunik-Yu_2025_BDD-Pooling establish their formal results leveraging concepts from geometric measure theory (federer2014geometric). After reviewing these ideas, we briefly discuss how estimation and inference can be conducted using the state-of-the-art methods from the RD design literature (Calonico-Cattaneo-Titiunik_2014_ECMA; Calonico-Cattaneo-Farrell-Titiunik_2019_RESTAT; Calonico-Cattaneo-Farrell_2020_ECTJ). Some of this work is conceptually related to a recent paper by chen2025semiparametric, who study estimation and inference for integral functionals on submanifolds when using nonparametric sieve estimation.
@@
Section 4 discusses heterogeneity analysis and subsequent aggregation of causal treatment effects along the boundary. Although the boundary average treatment effect parameter emerging from pooling-based methods discussed in Section 3 is often of interest, and provides a useful aggregate measure of average treatment effects, the richness of the two-dimensional assignment in BD designs offer the opportunity to explore heterogeneous treatment effects along the assignment boundary. In geographic applications, this heterogeneity is directly linked to the geographic location of the units, providing useful information about how treatment effects vary in space; in non-geographic settings, the heterogeneity is directly linked to features of the units that are captured by the bivariate score variable, and hence the interpretation is necessarily application-specific. Thus, Section 4 begins by introducing the boundary average treatment effect curve, which captures the average treatment effect for each point along the assignment boundary. We also discuss how other aggregate causal effects can be recovered from that building block parameter, such as the weighted boundary average treatment effect and the largest average treatment effect.
@@
After introducing the main ideas and causal parameter of interest, Section 4 reviews the theoretical and methodological takeaways from Cattaneo-Titiunik-Yu_2025_BDD-Distance and Cattaneo-Titiunik-Yu_2025_BDD-Location. These two recent papers investigate identification, estimation, and inference for heterogeneous treatment effects along the assignment boundary, and transformations thereof, from two connected but distinct perspectives: Cattaneo-Titiunik-Yu_2025_BDD-Distance studies boundary-point specific distance-based methods, which are directly motivated by the poloing-based approaches discussed in Section 3, while Cattaneo-Titiunik-Yu_2025_BDD-Location studies methods based on the bivariate location directly. For each approach, the estimation and inference results are valid both pointwise and uniformly along the assignment boundary, in addition to providing the basis for constructing other treatment effect estimators.
@@
Finally, Section 5 offers recommendations for practice, and concludes. The methods discussed in this chapter can be implemented using the general-purpose software packages rdrobust, rdhte, rdmulti, and rd2d, located at https://rdpackages.github.io/, where replication files and related references are also available.
@@ 2 The Boundary Discontinuity Design
Each unit $i$ in the study, $i=1,\ldots,n$, has a continuous bivariate score $\mathbf{X}_{i}=(X_{1i},X_{2i})$ that takes values in the set $\mathcal{X}\subseteq\mathbb{R}^{2}$. The assignment of units to the treatment or control condition depends on the location of their score $\mathbf{X}_{i}$ relative to a known one-dimensional boundary curve $\mathcal{B}$ that splits $\mathcal{X}$ into two disjoint regions: $\mathcal{X}=\mathcal{A}_{0}\cup\mathcal{A}_{1}$, with $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ the control and treatment disjoint (connected) regions, respectively, and $\mathcal{B}=\mathtt{bd}(\mathcal{A}_{0})\cap\mathtt{bd}(\mathcal{A}_{1})$, where $\mathtt{bd}(\mathcal{A}_{t})$ denotes the boundary of the set $\mathcal{A}_{t}$. (We follow convention and assume that the boundary belongs to the treatment group, that is, $\mathtt{bd}(\mathcal{A}_{1})\subset\mathcal{A}_{1}$ and $\mathcal{B}\cap\mathcal{A}_{0}=\emptyset$.) The observed outcome variable is $Y_{i}$.
@@
An ancestor of the BD design was used by Card-Krueger_1994_AER in their seminal study of the effects of the minimum wage on employment. In their analysis, the treatment of interest is the increase of the state minimum wage in New Jersey that went into effect on April 1, 1992, and rose the minimum wage to \$5.05. The authors compared a sample of fast food restaurants in New Jersey to a sample of fast food restaurants in eastern Pennsylvania, a state that shares a border with New Jersey and where the minimum wage did not increase. After the increase in New Jersey, the average starting wage was \$5.08 in the New Jersey sample and \$4.62 in the Pennsylvania sample.
@@
The rationale for comparing restaurants in New Jersey to restaurants in eastern Pennsylvania was that, by focusing on close geographic areas, these restaurants would be subject to similar economic conditions that would otherwise confound the effects of changes to the minimum wage. Although Card-Krueger_1994_AER also relied on a difference-in-differences design comparing both samples before and after the increase, they explicitly mentioned the close geographic proximity between the areas to justify their choice of comparison group, as they believed that fast-food stores in eastern Pennsylvania formed a “natural basis for comparison with the experiences of restaurants in New Jersey” due to Pennsylvania being a “nearby state[s]” (Card-Krueger_1994_AER, p. 773).
@@
The use of close proximity as the basis for reducing potential confounders and enhancing the credibility of causal interpretations bears a direct connection to the BD design. The main difference is that Card-Krueger_1994_AER included non-adjacent areas in New Jersey and Pennsylvania (e.g., the New Jersey sample included restaurants in the New Jersey shore, far away from the Pennsylvania border), and did not directly use the unit’s distance to the boundary in their analysis. The explicit use of distance to the assignment boundary is central to the BD design, which is why we consider Card-Krueger_1994_AER to be an ancestor rather than an instance of the BD design itself: while the authors leverage the change in treatment assignment induced by a geographic boundary, they did not explicitly localize to the boundary itself.
@@
holmes1998effect and black1999better are the earliest empirical examples that we could find employing the BD design. holmes1998effect used a BD design to study the effect of pro-business policies on manufacturing, comparing states that adopted right-to-work laws with adjacent states that did not. Using the longitude and latitude coordinates of each county’s centroid, he calculated the minimum distance of the centroid to the assignment border, and focused the analysis on counties close to the border as a way to “control for differences across states in these various characteristics that are unrelated to policy” (holmes1998effect, p. 671). Similarly, black1999better used a BD design to infer the quality of public schools from housing prices. She compared house prices on opposite sides of school attendance district boundaries, restricting the sample to houses close to the boundary. The justification for this localization was that houses on opposite sides of the boundary within a small area around it would be similar in all characteristics except for school quality—as the latter changes discontinuously at the attendance district boundary. In these examples, the bivariate score is the pair of latitude and longitude coordinates that determines the location of each unit (county or house), and the assignment boundary is the geographic border between two adjacent states or school attendance districts; the treatment is the policy or feature that changes abruptly at the border—labor policy for holmes1998effect and school quality for black1999better. This type of BD design is often referred to as a Geographic RD design.
@@
One of the most common non-geographic applications of the BD design occurs in education when a treatment is given on the basis of two exam scores. For example, ou2010leave studies the effect of a high school exit exam on the likelihood of dropping out of high school. The exam has a mathematics and a language arts component, each of which is graded separately; students must achieve a minimum proficiency in each section in order to pass the exam. The bivariate score in this case is the pair of mathematics and language test scores, and the treatment is passing the exam.
@@
Figure 1 illustrates the boundary and treatment assignment regions in two stylized settings corresponding to the geographic and non-geographic examples. Figure 1(a) illustrates the geographic BD design case, where units are split into adjacent treated and control areas according to their location with respect to a geographic border, as in black1999better and holmes1998effect. (This figure shows the New Jersey-Pennsylvania border as considered by Card-Krueger_1994_AER.) Figure 1(b) illustrates the non-geographic setting, using as example the study by ou2010leave where students receive two exam scores and a treatment is given only to those individuals who score above a minimum cutoff in each exam. In geographic applications, the bivariate score is always composed of the geographic coordinates of the units. In non-geographic applications, the bivariate score could include different types of variables; in addition to test scores from multiple exams, non-geographic bivariate scores include shares of airline passengers in origin and destination cities (snider2015barriers), tax rates and income ratios (egger2015impact), systolic and diastolic blood pressure measurements (dai2022effects), different component scores in a means testing formula (salti2022impact; kampfen2024heterogeneous), and population counts at different time periods (hinnerich2014democracy).
@@
Figure 1 illustrates an important distinction between the geographic and non-geographic examples. In non-geographic examples, the function that describes the boundary tends to be known, and is typically linear with few irregularities and very few kink points. In contrast, in geographic BD designs the boundary is typically the border between political or administrative subdivisions; as a consequence, its shape is only available as coordinates on a map, and its geometry can be complicated. In the best case possible, it will be piecewise linear with many kink points, but it will be more irregular in many applications. In his early study, holmes1998effect already recognized this challenge and discussed explicitly the complex boundary shapes created by geographic borders. Indeed, since the influential work by mandelbrot1967long; mandelbrot1983fractal, there has been an ongoing debate among mathematicians and philosophers trying to decide whether geographic borders (and other shapes in nature) are fractals (see avnir1998geometry, and references therein). From a methodological perspective, as recently formalized in Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location; Cattaneo-Titiunik-Yu_2025_BDD-Pooling, the geometry of the assignment boundary in BD designs has fundamental implications for identification, estimation, and inference. We will review some of these recent findings in the upcoming sections.
@@
In the standard RD design, the distance between the univariate score and the single cutoff is naturally always measured by the Euclidean distance, and observations are “close” to the cutoff $c$ when their univariate score $X_{i}$ implies a “small” distance $|X_{i}-c|$. In contrast, in the BD design, “closeness” to the boundary can be defined in different ways, depending on whether the parameter of interest is defined relative to a specific point on the boundary, or as some aggregate thereof along the entire boundary, such as an average or supremum. Furthermore, due its multivariate nature, the analysis can depend on the specific notion of distance used.
@@
In the next sections, we discuss the two main approaches to the analysis of BD designs. The first approach uses the variation in treatment assignment induced by entire boundary simultaneously, and thus pools all observations sufficiently “close” the boundary in a single analysis, producing a single average treatment effect across all boundary points. This pooling approach relies on localization to $\mathcal{B}$ by first computing the distance between each observation’s location and the nearest point on the boundary, denoted by $D_{i}$, and then retaining only those observations for which this distance is no larger than a specific bandwidth, denoted by $h$. The advantage of this approach is that it naturally mimics the univariate RD design, using the bivariate location of each unit to build a univariate score that is used directly in a “standard” univariate RD analysis. This approach thus produces a single average treatment effect estimate.
@@
In contrast, a second approach focuses on estimating the average treatment effect at each point on the boundary, performing localization relative to the specific point of analysis, thereby retaining observations whose distance to that point is no greater than a chosen bandwidth. Estimation and inference are conducted separately at each boundary point $\mathbf{x}\in\mathcal{B}$, which captures the heterogeneity of treatment effects along the boundary. Aggregate causal effects can then be recovered by averaging (or applying other transformations to) these boundary-point specific average treatment effects, including as a special case the average treatment effect along the entire boundary that is the focus of the pooling approach. The boundary-point-specific approach can be seen as focusing on heterogeneous average treatment effects that can then be used as building blocks to construct more aggregated causal parameters. The advantage of this approach is that it captures the full richness of the BD design by first learning the heterogeneity of the average treatment effect along the boundary, which usually summarizes unit characteristics of high policy relevance such as geographic location, academic performance, or economic need.

### v2

@@ Abstract
The boundary discontinuity (BD) design is a non-experimental method for identifying causal effects that exploits a thresholding rule based on a bivariate score and a boundary curve. This widely used method generalizes the univariate regression discontinuity design but introduces unique challenges arising from its multidimensional nature. We synthesize over $80$ empirical papers that use the BD design, tracing the method’s application from its formative stages to its implementation in modern research. We also overview ongoing theoretical and methodological research on identification, estimation, and inference for BD designs employing local polynomial regression, and offer recommendations for practice.
Keywords: boundary discontinuity, regression discontinuity, regression discontinuity with multiple scores, treatment effects estimation, causal inference
@@ 1 Introduction
The Regression Discontinuity (RD) design is one of the leading observational methods for program evaluation and causal inference (see Abadie and Cattaneo, 2018; Hernán and Robins, 2020, for overviews and further references). In its canonical form, a binary treatment is assigned to units whose value of a univariate score is equal to or above a known scalar cutoff, and not assigned to units whose score value is below the cutoff. Under the assumption that all observable and unobservable pretreatment characteristics vary smoothly at the cutoff, the discontinuous change generated by the hard-thresholding treatment assignment rule can be used to learn about causal treatment effects. Lee and Lemieux (2010) offer an early review and list of RD empirical applications, Cattaneo and Titiunik (2022) give a recent review of the RD methodological literature, and Cattaneo et al. (2020, 2024) provide a practical introduction to modern RD methods.
@@
The Boundary Discontinuity (BD) design generalizes the RD design: the score has two dimensions instead of one, and the treatment is assigned to units according to the location of their score relative to a known boundary curve that splits the support into two disjoint regions. This setup is sometimes called the Multi-Score RD design (Papay et al., 2011; Reardon and Robinson, 2012; Wong et al., 2013), or the Geographic RD design (Keele and Titiunik, 2015; Keele et al., 2015; Keele and Titiunik, 2016; Keele et al., 2017; Galiani et al., 2017; Rischard et al., 2021; Diaz and Zubizarreta, 2023). We discuss the empirical strategies most commonly employed for the analysis and interpretation of BD designs, and review ongoing theoretical and methodological results characterizing their econometric properties (Cattaneo et al., 2025a, b, 2026).
@@
We start in Section 2 by introducing the BD design and presenting a review of the empirical literature in Economics, Political Science, Education, and other disciplines. We report over $80$ empirical papers in Table 5, which form the basis of the rest of our discussion. We find that the overwhelming majority of papers collapse the bivariate score into a univariate distance score measuring the shortest distance from the unit’s location to the treatment assignment boundary, and then report a single average treatment effect estimated by pooling all the observations that are close to the boundary, regardless of their specific location. We refer to this empirical strategy as pooling-based methods. In contrast, only a handful of empirical papers have investigated heterogeneity by localizing to specific regions or points on the boundary, despite the usefulness of this rich information for policy evaluation and decision-making.
@@
In Section 3, we discuss pooling-based methods. We begin with the seminal work of Card and Krueger (1994). Although their paper is often cited as a canonical example of a difference-in-differences design, these authors are one of the first to argue that reducing geographic distance may increase comparability among units. We then discuss Holmes (1998) and Black (1999), who are among the first to employ localization to the assignment boundary explicitly as the central empirical strategy. Subsequent empirical work has recognized the importance of localization to the boundary, and has considered different empirical strategies for treatment effect estimation via local flexible regression methods. One of the most influential papers is Dell (2010), who incorporated a polynomial expansion of the bivariate location score as part of the local regression specification. More recently, an alternative empirical approach views the estimation based on distance to the boundary as a pooled univariate RD design, and thus estimates treatment effects using a local regression including a polynomial expansion of the univariate distance score—recent examples include Ito and Zhang (2020) and Dehdari and Gehring (2022). Many empirical papers also include boundary-segment fixed effects in their local regression specifications, and some include interactions with the treatment indicator and/or between the polynomial expansions of the univariate distance score and the bivariate location score. Furthermore, either to simplify the problem or because distinct treatments occur in different regions of the boundary, some papers reduce the BD design to boundary-segment-specific univariate RD designs—see Ou (2010), Londoño-Vélez et al. (2020) and Salti et al. (2022). We synthesize these various empirical strategies into eight distinct pooled local regression specifications.
@@
While pooling-based methods are widely used in empirical work, their econometric properties are not well understood. In ongoing research, Cattaneo et al. (2026) formally study identification, estimation, and inference for these methods, recast as a univariate RD design based on the shortest distance to the assignment boundary. The results leverage geometric measure theory (Federer, 2014) upon recognizing the assignment boundary as a one-dimensional submanifold curve on the plane, and enable estimation and inference using state-of-the-art RD methods (Calonico et al., 2014, 2019, 2020). Some of this work is conceptually related to a recent paper by Chen and Gao (2025), who study estimation and inference for integral functionals on submanifolds when using nonparametric sieve estimation.
@@
The boundary average treatment effect parameter emerging from the pooling-based methods is useful, but the richness of the two-dimensional BD assignment offers the opportunity to explore heterogeneous treatment effects along the assignment boundary. In geographic applications, the heterogeneity is directly linked to the geographic location of the units, providing useful information about how treatment effects vary in space; in non-geographic settings, the heterogeneity is directly linked to unit features that are captured by the bivariate score, and hence its interpretation is necessarily application-specific.
@@
In Section 4, we discuss heterogeneity analysis and subsequent aggregation of causal treatment effects in BD designs using the boundary average treatment effect curve (BATEC), which captures the average treatment effect for each point along the assignment boundary. We also summarize the results in Cattaneo et al. (2025a) and Cattaneo et al. (2025b). These papers study identification, estimation, and inference for BATEC and transformations thereof, from two distinct perspectives: Cattaneo et al. (2025a) studies boundary-point-specific distance-based methods, which are directly motivated by the pooling-based approaches discussed in Section 3, while Cattaneo et al. (2025b) studies methods based on the untransformed bivariate location. For each approach, the local polynomial estimation and inference results are valid both pointwise and uniformly along the assignment boundary, in addition to providing the basis for constructing other treatment effect estimators.
@@
We offer recommendations for practice and concluding remarks in Section 5. Our recommended methods can be implemented using the general-purpose software packages rdrobust, rdhte, rdmulti, and rd2d, located at https://rdpackages.github.io/, where readers can also find replication files and related references.
@@ 2 The Boundary Discontinuity Design
Each unit $i$ in the study, $i=1,\ldots,n$, has a continuous bivariate score $\mathbf{X}_{i}=(X_{1i},X_{2i})$ that takes values in the set $\mathcal{X}\subseteq\mathbb{R}^{2}$. The assignment of units to the treatment or control condition depends on the location of their score $\mathbf{X}_{i}$ relative to a known one-dimensional boundary curve $\mathcal{B}$ that splits $\mathcal{X}$ into two disjoint regions: $\mathcal{X}=\mathcal{A}_{0}\cup\mathcal{A}_{1}$, with $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ the disjoint (connected) control and treatment regions, respectively, and $\mathcal{B}=\mathtt{bd}(\mathcal{A}_{0})\cap\mathtt{bd}(\mathcal{A}_{1})$, where $\mathtt{bd}(\mathcal{A}_{t})$ denotes the boundary of the set $\mathcal{A}_{t}$. We assume that the boundary belongs to the treatment group, that is, $\mathtt{bd}(\mathcal{A}_{1})\subset\mathcal{A}_{1}$ and $\mathcal{B}\cap\mathcal{A}_{0}=\emptyset$. The observed outcome variable is $Y_{i}$.
@@
An ancestor of the BD design was used by Card and Krueger (1994) in their seminal study of the effects of the minimum wage on employment. In their analysis, the treatment of interest is the increase of the state minimum wage in New Jersey adopted on April 1, 1992, which rose the minimum wage to \$5.05. The authors compared a sample of fast food restaurants in New Jersey to a sample of fast food restaurants in eastern Pennsylvania, a state that shares a border with New Jersey and where the minimum wage did not increase. After the increase in New Jersey, the average starting wage was \$5.08 in the New Jersey sample and \$4.62 in the Pennsylvania sample.
@@
The rationale for comparing restaurants in New Jersey to restaurants in eastern Pennsylvania was that, by focusing on close geographic areas, these restaurants would be subject to similar economic conditions that would otherwise confound the effects of changes to the minimum wage. Although Card and Krueger’s main research strategy was a difference-in-differences strategy comparing NJ and PA before and after the increase, they explicitly mention the close geographic proximity between the areas to justify their choice of the comparison group, as they believed that fast-food stores in eastern Pennsylvania formed a “natural basis for comparison with the experiences of restaurants in New Jersey” due to Pennsylvania being a “nearby state” (Card and Krueger, 1994, p. 773).
@@
The use of close proximity as the basis for reducing potential confounders and enhancing the credibility of causal interpretations bears a direct connection to the BD design. The main difference is that Card and Krueger (1994) included non-adjacent areas in New Jersey and Pennsylvania (e.g., the New Jersey sample included restaurants in the New Jersey shore, far away from the Pennsylvania border), and did not directly use the unit’s distance to the boundary in their analysis. The explicit use of distance to the assignment boundary is central to the BD design, which is why we consider Card and Krueger (1994) to be an ancestor rather than an instance of the BD design itself: while the authors leverage the change in treatment assignment induced by a geographic boundary, they did not fully localize to the boundary itself in their analysis. (See Dube et al. (2010) for an example of a minimum wage study in the spirit of Card and Krueger (1994) that uses distance explicitly in the analysis.)
@@
Holmes (1998) and Black (1999) are the earliest empirical examples employing the BD design that we could find. Holmes (1998) used a BD design to study the effect of pro-business policies on manufacturing, comparing states that adopted right-to-work laws with adjacent states that did not. Using the longitude and latitude coordinates of each county’s centroid, he calculated the minimum distance of the centroid to the assignment border, and focused the analysis on counties close to the border as a way to “control for differences across states in these various characteristics that are unrelated to policy” (Holmes, 1998, p. 671). Similarly, Black (1999) used a BD design to infer the quality of public schools from housing prices. She compared house prices on opposite sides of school attendance district boundaries, restricting the sample to houses close to the boundary. The justification for this localization was that houses on opposite sides of the boundary within a small area around it would be similar in all characteristics except for school quality—as the latter changes discontinuously at the attendance district boundary. In these examples, the bivariate score is the pair of latitude and longitude coordinates that determines the location of each unit (county or house), and the assignment boundary is the geographic border between two adjacent states or school attendance districts; the treatment is the policy or feature that changes abruptly at the border—labor policy for Holmes (1998) and school quality for Black (1999). This type of BD design is often referred to as a Geographic RD design.
@@
One of the most common non-geographic applications of the BD design occurs in education when a treatment is given on the basis of two exam scores. For example, Ou (2010) studied the effect of a high school exit exam on the likelihood of dropping out of high school. The exam had a mathematics and a language arts component, each of which was graded separately; students must achieve a minimum proficiency in each section in order to pass the exam. The bivariate score in this case is the pair of mathematics and language test scores, and the treatment is passing the exam.
@@
Figure 1 illustrates the boundary and treatment assignment regions in two stylized settings corresponding to the geographic and non-geographic examples. Figure 1(a) illustrates the geographic BD design case, where units are split into adjacent treated and control areas according to their location with respect to a geographic border, as in Black (1999) and Holmes (1998). (This figure shows the New Jersey-Pennsylvania border.) Figure 1(b) illustrates the non-geographic setting, using as example the study by Ou (2010) where students receive two exam scores and a treatment is given only to those individuals who score above a minimum cutoff in each exam. In geographic applications, the bivariate score is always composed of the geographic coordinates of the units. In non-geographic applications, the bivariate score could include different types of variables; in addition to test scores from multiple exams, examples include shares of airline passengers in origin and destination cities (Snider and Williams, 2015), tax rates and income ratios (Egger and Wamser, 2015), systolic and diastolic blood pressure measurements (Dai et al., 2022), different component scores in a means-testing formula (Salti et al., 2022; Kämpfen and Mosca, 2024), and population counts at different time periods (Hinnerich and Pettersson-Lidbom, 2014).
@@
Figure 1 illustrates an important distinction between the geographic and non-geographic examples. In non-geographic examples, the function that describes the boundary tends to be known and is typically linear with few irregularities and few kink points. In contrast, in geographic BD designs the boundary is typically the border between political or administrative subdivisions; as a consequence, its shape is only available as coordinates on a map, and its geometry is complex: in the best case, it will be piecewise linear with kink points, but it will be more irregular in many applications. In his early study, Holmes (1998, pp. 681–682) recognized this challenge and discussed explicitly the complex boundary shapes created by geographic borders. Indeed, since the influential work by Mandelbrot (1967, 1983), there has been an ongoing debate among mathematicians and philosophers about whether geographic borders (and other shapes in nature) are fractals (see Avnir et al., 1998, and references therein). From a methodological perspective, as recently formalized in Cattaneo et al. (2025a, b, 2026), the geometry of the assignment boundary in BD designs has fundamental implications for identification, estimation, and inference. We will review some of these recent findings in the upcoming sections.
@@
In the standard RD design, the distance between the univariate score and the single cutoff is naturally measured by the Euclidean distance, and observations are close to the cutoff $c$ when their univariate score $X_{i}$ implies a small distance $|X_{i}-c|$. In contrast, in the BD design, closeness to the boundary can be defined in different ways, depending on whether the parameter of interest is defined relative to a specific point on the boundary, or as an aggregate thereof along the entire boundary, such as an average or a supremum. Furthermore, due to its multivariate nature, the analysis can depend on the specific notion of distance used.
@@
In the upcoming sections, we discuss the two main approaches to the analysis of BD designs. The first approach uses the variation in treatment assignment induced by the entire boundary simultaneously, and thus pools all observations sufficiently close to the boundary in a single analysis, producing a single average treatment effect across all boundary points. This pooling approach relies on localization to $\mathcal{B}$ by first computing the distance between each observation’s location and the nearest point on the boundary, denoted by $D_{i}$, and then retaining only those observations for which this distance is no larger than a specific bandwidth, denoted by $h$. The advantage of this approach is that it naturally mimics the univariate RD design, using the bivariate location of each unit to build a univariate score that is used directly in a standard univariate RD analysis. This approach produces a single average treatment effect estimate.
@@
The second approach focuses on estimating the average treatment effect at each point on the boundary, performing localization relative to the specific point of analysis, thereby retaining observations whose distance to that point is no greater than a chosen bandwidth. Estimation and inference are conducted separately at each boundary point $\mathbf{x}\in\mathcal{B}$, capturing the heterogeneity of treatment effects along the boundary. Aggregate causal effects can be recovered by averaging (or applying other transformations to) these boundary-point average treatment effects, including as a special case the average treatment effect along the entire boundary that is the focus of the pooling approach. The boundary-point approach can be seen as focusing on heterogeneous average treatment effects that are then used as building blocks to construct more aggregate causal parameters. The advantage of this approach is that it captures the full richness of the BD design by first learning the heterogeneity of the average treatment effect along the boundary, which usually summarizes relevant unit characteristics such as geographic location, academic performance, or economic need.